### v1

# OCCAM: Online Continuous Controller Adaptation with Meta-Learned Models

Abstract
Control tuning and adaptation present a significant challenge to the usage of robots in diverse environments. It is often nontrivial to find a single set of control parameters by hand that work well across the broad array of environments and conditions that a robot might encounter. Automated adaptation approaches must utilize prior knowledge about the system while adapting to significant domain shifts to find new control parameters quickly. In this work, we present a general framework for online controller adaptation that deals with these challenges. We combine meta-learning with Bayesian recursive estimation to learn prior predictive models of system performance that quickly adapt to online data, even when there is significant domain shift. These predictive models can be used as cost functions within efficient sampling-based optimization routines to find new control parameters online that maximize system performance. Our framework is powerful and flexible enough to adapt controllers for four diverse systems: a simulated race car, a simulated quadrupedal robot, and a simulated and physical quadrotor.
Keywords: Controller Adaptation, Robot Model Learning, Meta-Learning
1 Introduction
Figure 1: We demonstrate a flexible method for controller optimization based on online adaptation of meta-learned models on diverse robots.
Robust and highly-performant control policies are critical for the successful application of robots in diverse environments. All control algorithms for robots have parameters that must be carefully tuned for good performance, but whose optimal values are generally not obvious a priori and are not easily computed. In practice, these parameters are often tuned based on the designer’s intuition and using heuristics that do not align with the actual downstream task of the robot–for instance, gain tuning a quadrotor using the step response. This problem is often exacerbated by the sim-to-real gap [1], requiring designers to fine-tune control parameters directly on the real system. Furthermore, the optimal control parameters can vary over time if the robot’s physical attributes change or the robot enters new environments. Modern data-driven approaches that adapt to these changes usually perform online system identification (potentially in a latent space) for optimization-based control or policy learning [2, 3, 4, 5]. These adaptive methods have shown strong results, though for many of these approaches, it is difficult to tune their behaviors after training, and their generalizability to novel domains is unclear.
Our main contribution is OCCAM, a framework for online controller adaptation to new environments using adaptable learned models of performance metrics. We combine meta-learning with Bayesian regression to train predictive models on simulated data that provide a strong prior on controller performance and are readily fit to data gathered during operation. These models are then used as cost functions to search for optimal control parameters in new environments unseen during training. One of the major advantages is the flexibility of our framework: it can easily be applied to any system with a parameterized controller. To illustrate this, we demonstrate that our method performs online controller adaptation on a simulated race car equipped with a nonlinear control law, a simulated and physical quadrotor micro-aerial-vehicle (MAV) with a model-based geometric controller, and a simulated quadruped with a neural-network locomotion policy.
Figure 2: An overview of our method and predictive model $\hat{f}$. Given previous sensor measurements and inputs from the system, the Optimization phase uses the prediction model to search for the best gain to try next. This gain is sent to the real system for the Evaluation phase, during which sensor measurements and performance measures are collected. The Adaptation phase computes new weights to update the prediction model with the information gathered during the evaluation. Arrows in red indicate predicted and estimated quantities, while arrows in black indicate signals from the actual system.
2 Related Work
Automated Controller Tuning: Approaches for controller tuning often use Bayesian Optimization (BO) with Gaussian Process (GP) surrogate models [6, 7, 8, 9], probabilistic sampling-based approaches [10, 11, 12], or gradient descent through the controller and dynamics [13]. Prior data can be incorporated into BO by simply adding datapoints to the GP or specifying a prior mean function. [14, 15] add prior datapoints to a GP to find control policies for damaged robots online; [16, 17] instead use offline data to learn fixed kernel functions that make BO more efficient. These methods can degrade if the kernel or prior becomes inaccurate as the true underlying function changes. Our method is instead trained to accommodate these domain shifts.
Adaptation in Data-Driven Control: A closely-related line of works combines Reinforcement Learning (RL) for learning low-level control policies with system identification. Short-term histories of states and actions are encoded as latent variables to condition either model-free policies [5, 18, 19, 20, 21] or dynamics models [22], allowing the model to modulate its behavior in different environments. [23, 24, 25, 26] instead learn a repertoire of diverse policies for varying environments. The deep context encoders common in these works might not be accurate when the test system lies outside training distribution. A key idea in our method is to explicitly adapt the model to the observed measurements online.
Meta-Learning for Robotics: Meta-learning aims to deal with the lack of adaptability of learned models by incorporating the adaptation procedure into the training loss to train models that are well-suited to efficiently fit new data. [27, 28, 29] use meta-learning to train adaptable mean, kernel, and acquisition functions for BO with applications to controller tuning. [30] use Bayesian regression to meta-learn dynamics models that can be adapted to new dynamics at test-time. [21, 31, 32, 33] employ meta-learning to train low-level control policies with RL that can be adapted quickly to new contexts. In this work, we alleviate overfitting issues common to these meta-learning methods [34] by adapting our model in a low-dimensional space and training it to characterize the uncertainty in its predictions.
Adaptive Control: Adaptive control generally seeks to augment the control inputs to a system to cancel out disturbances and align the system’s closed-loop dynamics to those of a desired reference model [35]. Direct approaches such as $\mathcal{L}$1-Adaptive control and Model-reference adaptive control have successfully been applied to robotic aerial vehicles [36, 37, 38, 39], manipulators [40], and bipedal robots [41]. Indirect approaches based on online model learning [2, 3, 42] attempt to estimate the dynamics of the robot in order to compute the control actions. Poor values of the additional tunable adaptive gains in adaptive control can render the controller unstable [43]. It is also unclear how a reference model that is always feasible for the actual system to track might be synthesized, and adaptive control generally requires strong assumptions about model structure and differentiability [38, 39], which our method does not require.
3 Problem Statement
In this work we address the problem of continuously adapting a controller using a combination of offline simulation data and data gathered from the robot online. The goal of the OCCAM procedure is then to find the control gains at iteration $k$, $g_{k}$, that maximize the reward over some time horizon $\Delta T$:
$g^{*}$ $=\operatorname*{arg\,max}_{g}J(y_{k+\Delta T})=\operatorname*{arg\,max}_{g}J(f_{\theta_{k}}(z_{k:k+\Delta T},u_{k:k+\Delta T},\tau_{k}))$ (1)
Where we let $z_{k}\in\mathbb{R}^{N_{z}}$ denote a vector containing the sensor observations from the system at the $k$’th discrete timestep and $u_{k}\in\mathbb{R}^{N_{u}}$ denote the control inputs applied to the system by the controller. $y_{k+\Delta T}\in\mathbb{R}^{N_{y}}$ denotes a vector of performance measures, such as tracking error, control effort, or speed, computed as a fixed function of the measurements and control history over the horizon $\Delta T$: $y_{k+\Delta T}=f(z_{k:k+\Delta T},u_{k:k+\Delta T},\tau_{k})$. In addition, we specify a reward function, $J$ that maps the vector of performance measures onto a single scalar value: $J(y)\in\mathbb{R}$.
The dynamics of our system are governed by a set of unobserved intrinsic and extrinsic system parameters such as inertial parameters or friction coefficients. We refer to these system parameters collectively as $\theta_{k}$. When generating data in simulation the designer would typically perform domain randomization by sampling values from some subset $\Theta$ to explore the behavior of the system in different regimes.
We are also given a control policy, $C$ (e.g. PID, MPC, etc.) that computes control inputs from previous measurements: $u_{k+1}=C_{g_{k}}(\tau_{k},z_{0:k},u_{0:k})$. $C$ is parameterized by tunable gains $g_{k}\in\mathbb{R}^{N_{g}}$, and $\tau_{k}$, which is the control objective of the controller derived from the robot’s task. For example, $\tau_{k}$ could denote an encoding of the upcoming trajectory that we want a mobile robot to follow.
4 Method
4.1 Modeling Controller Performance
The foundation of OCCAM is a predictive model $\hat{f}$ which is designed to predict future performance measures as a function of control parameters. The structure of this prediction model is shown in Figure 2 (left). The recent measurements and control inputs (from a sliding window of size $H$), $z_{k-H:k},u_{k-H:k}$, are fed into a shallow fully connected network that produces a context encoding [5] that captures relevant, observable aspects of the system. This encoded result is combined with an encoding of the current task, $\tau_{k}$ and the proposed control parameters, $g_{k}$ and fed into another fully connected network.
To quickly adapt to new data, our prediction network employs a last linear layer structure [44, 30] wherein the fully connected network produces a matrix $\Phi\in\mathbb{R}^{N_{y}\times N_{b}}$ which is combined with a weight vector $w_{t}\in\mathbb{R}^{N_{b}}$ to produce the final prediction: $\hat{y}_{t}=\Phi w_{t}$. We can think of the columns of the matrix $\Phi=[\phi_{1},...\phi_{N_{b}}]$ as a set of basis functions weighted by $w_{t}$. Here we use $t$ as an index instead of $k$ to denote that the collection of performance measures and adaptation occur at a different rate than the low-level control and sensing. We treat $w_{t}$ as a normally distributed random variable: $w_{t}\sim\mathcal{N}(\mu_{t},\Sigma_{t})$. Because $\hat{y}$ is a linear transformation of $w$, we obtain the following distribution for it: $\hat{y}_{t}\sim\mathcal{N}(\Phi\mu_{t},\;\Phi\Sigma_{t}\Phi^{T})$. We use a Kalman filter with identity dynamics to recursively estimate $w$ as new data, $(\Phi_{t},y_{t})$, is obtained online from the system:
$\bar{\Sigma}_{t+1}$ $=\Sigma_{t}+Q$ (2)
$K$ $=\bar{\Sigma}_{t+1}\Phi\left(\Phi\bar{\Sigma}_{t+1}\Phi^{T}+R\right)^{-1}$
$w_{t+1}$ $=w_{t}+K(y_{t+1}-\Phi w_{t})$
$\Sigma_{t+1}$ $=\left(I-K\Phi\right)\bar{\Sigma}_{t+1}$
Informally adopting standard Kalman Filter terminology, $Q$ is a positive definite matrix that scales with our confidence in keeping $w$ constant, while $R$ captures the error in our model, $\Phi$, and the noise in the measurements of $y$.
4.2 Meta-Training
Input: Training dataset $\mathcal{D}$, network parameters $\psi$, $w_{\text{pre}}$
Output: Trained parameters $\psi$, $w_{0},\Sigma_{0},Q,R$
Initialize $w_{0}=w_{\text{pre}}$ ;
Initialize $\Sigma_{0},Q,R$ = $I$ ;
foreach training epoch do
       foreach parameter dataset $D_{\theta}=(G,Z,U,\tau,Y)$in $\mathcal{D}$ do
             Sample random-size subset $\mathcal{D}_{\text{train}}$ of size $n$ from $\mathcal{D}_{\theta}$;
             Initialize $w_{0}^{\prime}=w_{0},\Sigma_{0}^{\prime}=\Sigma_{0}$ ;
             foreach datapoint $(g_{i},z_{i},u_{i},\tau_{i},y_{i})$ in $\mathcal{D}_{\text{train}}$ do
                   Compute $w_{i}^{\prime},\Sigma_{i}^{\prime}$ using (2)
             end foreach
            Compute $\mathcal{L}=\lVert\Phi_{\psi}(G,Z,U,\tau)w_{n}^{\prime}-Y\rVert^{2}$ ;
             foreach $p\in\{\psi,w_{0},\Sigma_{0},Q,R\}$ do
                   $p\leftarrow p-\alpha\nabla_{p}\mathcal{L}$ ;
             end foreach
       end foreach
end foreach
Algorithm 1 Meta-Learning with Kalman Filter Base Learner
Our goal is to produce learned basis functions that capture the variation we observe across different system parameters, $\theta$. Note that our approach does not attempt to enumerate or estimate $\theta$ directly, but instead adapts in the latent space of $w$. To achieve this, after pretraining our network for a few epochs using standard gradient descent, we switch to gradient-based meta-learning [31], which consists of a base learner which adapts the base model to data from each “task” (in our case, from systems with varying $\theta$), and an outer loop which adjusts the parameters of the base model based on the base learner’s loss. Our base learner is the recursive update (2) applied to a subset of each system dataset. Because every computation in the Kalman Filter is fully differentiable, once the final weights are computed, the gradient of the loss can be backpropagated through the full computation graph to update the neural network parameters and the Kalman filter parameters $w_{0},\Sigma_{0},Q,R$ using standard automatic differentiation tools such as PyTorch’s autograd engine [45]. Algorithm 1 shows how we use the standard meta-learning loops with our Kalman Filter base learner as the inner loop. Meta-learning $w_{0}$ and $\Sigma_{0}$ is similar in philosophy to Harrison et al. [30] and Bertinetto et al. [46], which also learn the prior distribution for the last-layer weights. However, they do not consider a Kalman Filter formulation for regression, and thus do not include process and measurement covariance matrices $Q$ and $R$.
4.3 Controller Optimization With Learned Model
We want to use the learned prediction model to solve (1) while also using our knowledge of the uncertainty on the model’s predictions. More specifically, at regular intervals OCCAM seeks to find a choice of control parameters that maximizes the uncertainty-aware reward:
$g^{*}$ $=\operatorname*{arg\,max}_{g}\mu(J(\hat{y}_{k+\Delta T}))+\beta\sigma(J(\hat{y}_{k+\Delta T}))$
where $\hat{y}_{k+\Delta T}=\Phi(g_{k},\tau,z,u)w_{k}$ is the output of our prediction model, and $\beta$ is a scalar penalty that controls the acceptable confidence bound. While many optimizers could be used, we employ a simple evolutionary modification to random-search in our evaluation settings. In each search iteration, we generate a fixed number of random samples from a uniform distribution $G_{r}=\mathcal{U}(g_{\text{min}},g_{\text{max}})$. We also apply $N_{r}$ random perturbations to the optimal gain vectors found in the $T$ previous trials, generating $T\times N_{r}$ evolutionary search samples $G_{e}$. The full set of samples is thus $G=G_{r}\cup G_{e}$. Each sample is concatenated with the upcoming command, $\tau_{k}$ and system history, $z_{k-H:k},u_{k-H:k}$, and passed through the network to obtain the expected reward. After evaluating all samples, we pick the sample with the highest expected reward.
5 Experimental Procedures
5.1 Simulated Evaluation Platforms
To show the flexibility of our framework, we evaluate it on three distinct simulated and one real robotic platform. In the following subsections, we give details on each system and explain how our method is applied. Extensive descriptions of each system and its accompanying controllers are provided in the supplementary material.
2D Race Car with Nonlinear Controller: Our first simulated robotic system is a 2-dimensional car racing around procedurally generated race tracks, modified from OpenAI’s gym environment [47]. The car has unknown mass, engine power, and tire friction as $\theta$, with an encoding of the race track shape as task $\tau$. OCCAM’s goal is to continually optimize the six-dimensional control gains $g$ of a PD steering controller that keeps the car on the centerline of the track, and a proportional acceleration controller ($C$) which contain speed limit parameters that enter nondifferentiably into the controller to set thresholds on the maximum allowable speed of the car in straights and corners. The reward function is a linear combination of the performance measures: inverse average lateral tracking error, inverse average wheelslip, and average speed over the upcoming track segment. OCCAM adapts and selects new gains every time the car traverses $\frac{1}{3}$ of a lap.
Quadrotor with Geometric Controller: Our second simulated platform is a quadrotor MAV equipped with a geometric trajectory tracking controller defined on SE(3) [48], implemented in RotorPy [49]. The quadrotor has five unknown system parameters $\theta$ which are the quadrotor’s true mass, principal moments of inertia, and thrust coefficient. The controller computes a feedforward motor speed command based on reference trajectory $\tau$ using the quadrotor’s nominal parameters (centered around those of the Crazyflie platform [50]) and a feedback command to correct tracking errors. The controller is parameterized by eight PD gains on the position and attitude. The reward function is a linear combination of the inverse average positional tracking error, inverse average yaw tracking error, inverse average pitch and roll, and inverse average commanded thrust over the episode. The adaptation frequency $\Delta T$ is 4 seconds. We also evaluate our framework on a physical Crazyflie with the same controller and performance measures. Details of the physical experiments are described in Section 6.2.
Quadrupedal Robot with Learned Locomotion Policy: Our third simulated robotic platform is a quadrupedal robot equipped with a static pretrained locomotion policy $\pi$ trained using model-free RL [26]. $\pi$ outputs joint angles at 50Hz such that the torso of the robot follows a velocity twist command $\tau_{k}=(\dot{x}_{\text{des}},\dot{y}_{\text{des}},\dot{\omega}_{\text{des}})$. We treat $\pi$ as our controller $C$ for this system. Although $\pi$ is parameterized by a deep neural network, it is also conditioned on an additional, eight-dimensional command $g_{k}$ that allows the user to specify high-level desired behaviors that the policy should follow: stepping frequency, body height, footswing height, stance width, and three discrete variables which jointly specify the quadrupedal gait. We treat $g_{k}$ as the controller parameters to be tuned automatically based on the quadruped’s randomized parameters and the twist command. The unknown system parameters $\theta_{k}$ for the quadruped are added mass payloads to the robot base, motor strengths, and the friction and restitution coefficients of the terrain. The adaptation frequency $\Delta T$ is 3 seconds.
5.2 Model Training and Testing
To generate training data from each simulated platform, we randomly sample $N$ parameter vectors $\theta_{0:N}$. For each $\theta_{i}$, we sample $N_{B}$ random $g$ and $\tau$, drive the system to a random initial dynamical state to collect history $z,u$, and finally roll-out the system with controller $g$ to gather performance metrics $y$. We train a separate basis-function network for each system.
We test each method on robots with system parameters sampled from outside of the range of the training dataset, testing the ability of each method to extrapolate outside of the training data. We evaluate methods on 15 out-of-distribution system parameters on multiple different tasks (racetracks for the racecar, ellipsoidal trajectories for the quadrotor, and twist commands for the quadruped) and 8 random seeds. For each platform, we fix the reward functions $J$ and uncertainty penalty $\beta$. Each test consists of multiple iterations of the “Optimize, Evaluate, Adapt” procedure shown in Figure 2. Because we run each test in an online setting with no resets between timesteps, if a method selects dangerous gains that cause a crash, the reward for every subsequent timestep of that evaluation is set to zero. More details on dataset sizes, model architecture, and training and testing parameter ranges are provided in the supplementary material.
5.3 Baselines
We compare our method on each test system against three baselines. The first baseline is Reptile [51], a first-order version of the common gradient-based meta-learning method MAML [31]. Reptile uses minibatched gradient descent as the base learner and it only outputs a single point prediction of $y$. For our second baseline, we compare against a GP model whose kernel function is the composition of a network trained only using phase-1 of our method with an RBF kernel, mirroring [17, 29]. We also compare against constant control parameters that were hand-tuned by experts for robust and conservative performance across a variety of commands and system parameters. In our quadrotor example, we additionally consider an $\mathcal{L}$1-Adaptive control [36] baseline adapted from [37], which augments the nominal control with an additional term based on an estimated disturbance vector.
To evaluate the usefulness of adapting the prediction model, we perform an ablation where we do not adapt $w$ from trial to trial and only use $w_{0}$ and $\Sigma_{0}$ in each trial (context-only). In the robotic experiments, the context-only ablation mimics methods that directly estimate latent system parameters from states and actions, such as Kumar et al. [5], since the prediction network only uses the history context encoder and does not utilize our last-layer weight adaptation. To evaluate the usefulness of meta-training, we also run an ablation where last-layer weight adaptation occurs, but the network is not trained using the meta-learning procedure (no-meta).
6 Results
Table 1: Average Final Reward and Crash Rate on Robotic Systems
Race Car Quadrotor Quadruped Method Avg Final Rwd ($\uparrow$) Crash % ($\downarrow$) Avg Final Rwd ($\uparrow$) Crash % ($\downarrow$) Avg Final Rwd ($\uparrow$) Crash % ($\downarrow$) Nominal $0.25\pm 0$ 8 $0.72\pm 0.24$ 33.8 $0.69\pm 0.08$ 8.8 LK-GP $0.29\pm 0.14$ 24.2 $1.11\pm 0.34$ 29.9 $0.74\pm 0.07$ 6.6 Reptile [51] $0.29\pm 0.25$ 14.2 $1.19\pm 0.37$ 33.8 $0.70\pm 0.09$ 8.1 $\mathcal{L}$1-Adaptive [36] - - $0.61\pm 0.25$ 33.6 - - OCCAM (no-meta) $0.1\pm 0.16$ 45.9 $1.05\pm 0.42$ 32.8 $0.73\pm 0.09$ 8.1 OCCAM (context-only) $0.22\pm 0.16$ 37.3 $1.08\pm 0.38$ 27.3 $0.73\pm 0.07$ 6.7 OCCAM (Ours) $\mathbf{0.40\pm 0.17}$ 13.4 $\mathbf{1.40\pm 0.30}$ 26.3 $\mathbf{0.75\pm 0.07}$ 6.3
(a) Plots of performance reward value vs time on robotic systems (Higher is Better)
Figure 3: Time vs. Reward curves on all test systems. Our method shows robust prior performance and adaptation across all tests. For the robotic systems, normalized reward is computed by standardizing rewards according to the training dataset statistics for that system and subsequently subtracting the nominal controller’s reward.
In Table 1 we report the crash rate and final average reward achieved by each method. To show the advantage of each method over simply using the expert-tuned nominal gains, in Figure 3 we plot the difference between each method’s reward and the nominal gain’s reward over time in each test. We report the raw performance metrics achieved by each method in the supplementary material.
OCCAM outperforms both ablations and all baselines on all benchmarks, demonstrating the utility of both meta-learning basis functions and weight adaptation. Adaptation with our method also occurs within a few timesteps, which corresponds to 10-20 seconds of data on each system. The quadruped environment is challenging in particular, as OCCAM must predict the performance of an expressive RL policy; nonetheless, it finds behavior parameters that improve the policy’s performance. The quadrotor also presents unique challenges, as the test parameters can have up to 300% parametric error from the nominal model that must be compensated by tuning PD feedback gains (the controller does not have an integral term). These extreme parametric errors also cause instability and poor tracking in the $\mathcal{L}$1-adaptive controller. Meanwhile, our method finds gains that result in the lowest crash rate and nearly a 50% reduction in tracking error compared to the nominal gains and $\mathcal{L}$1-Adaptive controller. Both Reptile and LK-GP saturate early in the evaluations. Reptile tends to overfit to the training set and online training data [34] and gets stuck in a loop of selecting the same gains and repeatedly fitting to them.
In the supplementary material, we additionally show that OCCAM also performs well on common global function optimization benchmarks, makes interpretable optimizations to the gains as system parameters change, and learns a structured latent weight space.
6.1 Weight Adaptation Enables Out-of-Distribution Generalization
To further demonstrate the necessity of model adaptation for controller tuning in out-of-distribution scenarios, we also evaluate each baseline on test scenarios that lie within the training distribution, and present a table of these results in the supplement. In these test cases, the context-only baseline performs comparably to our full method. Therefore, its performance drop in the out-of-distribution scenarios suggests that the learned context encoder alone is not sufficient to achieve generalization. Similarly, our method outperforms LK-GP because the usefulness of the latent space of the kernel network degrades as we move out of its training distribution.
To emphasize OCCAM’s generalizability, we also use OCCAM to tune the controller gains of a simulated Crazyflie to minimize tracking error in three different wind conditions up to 0.5 m/s, when wind was not encountered at all during training. OCCAM is able to achieve an average tracking error of under 7cm, outperforming all baselines including $\mathcal{L}$1-Adaptive control (10cm), shown in the supplementary material. Meanwhile, context-only achieves the worst tracking performance (30cm).
6.2 OCCAM crosses the Sim-to-Real Gap
Figure 4: Positional tracking error results on physical Crazyflie quadrotor following a 3-dimensional ellipsoidal reference trajectory.
Finally, we apply OCCAM to tune the gains of a physical Crazyflie quadrotor. The controller for the physical Crazyflie runs on a Robot Operating System (ROS) node running on a laptop with no GPU and a 1.6Ghz Intel i5 processor. The ROS node receives the quadrotor’s position through a Vicon motion-capture system [^1] https://www.vicon.com/ running at 100Hz and computes a desired collective thrust and quaternion attitude command for the quadrotor, which the onboard PID controller then converts into motor speeds. We train our predictive model on simulation data sampled from both the train and test parameter ranges, and add OCCAM as a lightweight layer that runs every 3 seconds on top of the ROS control node.
Position tracking error results for a three-dimensional ellipsoidal reference trajectory are shown in Figure 4. Control parameters tuned by OCCAM result in lower tracking error than the expert-tuned nominal controller and achieve comparable tracking error to the $\mathcal{L}$1-Adaptive controller. We also run additional experiments where a 5 gram mass is added to the Crazyflie (see Figure 1), which represents a 20% increase from the its nominal mass. Our method is able to find gains that decrease the quadrotor’s tracking error in the z-axis by 54% compared to the nominal controller and by 17% compared to the $\mathcal{L}$1-Adaptive controller. Figures of tracking errors from these physical Crazyflie experiments are shown in the supplementary material.
7 Discussion and Limitations
Our results demonstrate a single algorithm that optimizes controllers on a diverse array of robotic platforms using simple search techniques. Importantly OCCAM obviates the need for manual gain tuning that typically requires domain expertise, and instead provides automatic, continual controller adaptation. This allows designers to focus on choosing performance metrics better aligned with the actual task. OCCAM is also compatible with any controller with tunable parameters, allowing straightforward application to controllers that are nonlinear, discontinuous, and not analytic. As shown in our Crazyflie experiment, OCCAM’s models and optimization routine are lightweight enough to run in realtime without a GPU.
In this work, we tested OCCAM on commonly used controllers with relatively low-dimensional gain spaces, though one might also wish to tune a higher-dimensional policy, such as a neural network. Although the policy could be conditioned on a low-dimensional behavior vector, one could also treat the weights of either the entire neural network or its last few layers [52, 53] as the controller gains and optimize those using a more powerful optimizer such as gradient descent.
Beyond this, there are a few areas in which our method could be improved. If the stochastic optimizer makes a poor initial choice and the uncertainty penalty is poorly tuned, OCCAM sometimes settles early in a suboptimal local maximum. For some systems, this incorrect choice of gains can cause crashes. Incorporating a more sophisticated optimizer and exploration strategy would be interesting future work. Secondly, if the designer has knowledge of the true system parameters at test time, there is no clear way to set the weights $w$ besides through the training data distribution. Lastly, in this work we do not rigorously evaluate how far the generalizability of our method applies, or changing the reward function $J$ online. Future work includes addressing these limitations, along with demonstrating our method on broader classes of systems.